x + 2y = 1
7x - 5y = 26
I need to solve using substitution,

Answer by TimothyLamb(4379)

---
x + 2y = 1
7x - 5y = 26
---
as a check on the substitution method, let's first solve the given system of linear equations by using matrices:
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 3
y= -1
NOTE: the above solver uses the matrix method to solve the system.
---
now let's solve by substitution:
---
x + 2y = 1
x = 1 - 2y
---
7x - 5y = 26
7(1 - 2y) - 5y = 26
7 - 14y - 5y = 26
7 - 19y = 26
19y = -19
y = -1
---
x = 1 - 2y
x = 1 - 2(-1)
x = 1 + 2
x = 3
---
answer:
x= 3
y= -1
---
the matrix method and the substitution method agree, so we must have the correct answer !
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php